# Asymptotic independence in higher dimensions and its implications on risk management

### v2

@@ 3 Mutual asymptotic independence
Pairwise asymptotic independence has often either been used as a natural extension of asymptotic independence ([2, 18]), or, taken as a consequence from other relevant properties ([13, Remark 6.2.5]), or, implicitly assumed ([26]) in a variety of works. Next, we define a notion that captures the global joint concurrent tail behavior of random vectors portrayed by many popular multivariate dependence structures, e.g., dependence defined using Gaussian, Marshall-Olkin, or various Archimedean copulas, etc., but not restricted to mere replication of pairwise comparisons of tails.
@@
A random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with survival copula $\widehat{C}$ is mutually asymptotically independent if for all $S\subseteq\mathbb{I}_{d}$ with $|S|\geq 2$, we have
@@
When $d=2$, both (2.2) and (3.1) boil down to (2.1) and hence, are equivalent. Assuming $d\geq 3$ and mutual asymptotic independence, if we take all choices of $S\subseteq\mathbb{I}_{d}$ with $|S|=2$, then (3.1) is just a restatement of (2.2), implying pairwise asymptotic independence. We find that mutual asymptotic independence implies pairwise asymptotic independence, which we summarize next.
@@
If a random vector $\boldsymbol{Z}\in\mathbb{R}^{d},$ $d\geq 2$ is mutual asymptotically independent, then it is also pairwise asymptotically independent.
The reverse implication of (3.2) is not necessarily true as we see in the following example, which mimics the consequences for the analogous notions of classical mutual and pairwise independence ([22]).
@@
The difference between pairwise and mutual independence can be shown using an $\mathbb{R}^{3}$-valued random vector with Bernoulli marginals (cf. [22, Chapter 2]). We take a similar spirit using uniform marginals. Consider i.i.d. uniform [0,1] random distributed random variables $U,V,W$. Then $\boldsymbol{Z}=(U,V,W)$ is mutually asymptotically independent (cf. Example 3.4) and hence, pairwise asymptotically independent as well. Now consider $\boldsymbol{Z}=(Z_{1},Z_{2},Z_{3})$ such that
$\boldsymbol{Z}$ =$=\begin{cases}(U,V,\min(U,V)),&\text{ with prob. }\;\;1/3,\\ (U,\min(U,V),V),&\text{ with prob. }\;\;1/3,\\ (\min(U,V),U,V),&\text{ with prob. }\;\;1/3.\end{cases}$
First note that for $10<z<$ 1,
@@
and hence, the $Z_{j}$’s are identically distributed. If $\widehat{C}$ denotes the survival copula of $\boldsymbol{Z}$, then we can check that for any $\{j,\ell\}\subset\{1,2,3\}$,
$\widehat{C}_{j,\ell}(u,u)$ $=\mathbb{P}(Z_{j}>2-\sqrt{1+3u},Z_{\ell}>2-\sqrt{1+3u})$
@@
$=(\sqrt{1+3u}-1)^{2}=9u^{2}/4+o(u^{2}),\quad u\downarrow 0.$
Hence, $\boldsymbol{Z}$ exhibits pairwise asymptotic independence. But
@@ 3.1 Examples: Part I
In Proposition 2.6 we showed that any random vector $\boldsymbol{Z}$ with dependence given by the generalized Marshall-Olkin survival copula $\widehat{C}^{\mathrm{MO}(\Lambda)}$ as defined in (2.5) is pairwise asymptotically independent. The next proposition shows that it is indeed mutually asymptotically independent.
Proposition 3.6.
Let the dependence structure of a random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ be given by the generalized Marshall-Olkin survival copula $\widehat{C}^{\mathrm{MO}(\Lambda)}$ as in (2.5) with any choice of parameter set $\Lambda$. Then $\boldsymbol{Z}$ possesses mutual asymptotic independence.
Note that, in light of Proposition 3.2, clearly Proposition 2.6 on pairwise asymptotic independence is a direct consequence of Proposition 3.6.
@@ 3.2 Examples: Part II
In this section, we discuss examples that are pairwise asymptotically independent but sometimes are not mutually asymptotically independent. This will include a large class of examples from the Archimedean copula and as well the Gaussian copula family.
3.2.1 Archimedean copula.
@@
Let the dependence structure of a random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ be given by an Archimedean copula $C^{\phi}$ with generator $\phi$ as in (2.8). Suppose $\phi^{\leftarrow}$ is $d$-times continuously differentiable and $(-D)^{j}\phi^{\leftarrow}(0)<\infty$ $\forall\,j\in\mathbb{I}_{d}$. Then $\boldsymbol{Z}$ possesses both pairwise and mutual asymptotic independence.
The proof follows directly from Charpentier and Segers [4, Theorem 4.3]. The Archimedean copulas of Theorem 3.7 have the property that for any subset $S\subset\mathbb{I}_{d}$ with $|S|\geq 2$, the survival copula $\widehat{C}_{S}^{\phi}$ on the $|S|$-dimensional marginal behaves like the independence copula near the tails, i.e.,
@@
where the upper tail order of $C_{S}$ is $\kappa_{S}=|S|$ (also follows from [4, Theorem 4.3]). In particular, the upper tail order for $C^{\phi}$ is $\kappa=\kappa_{\mathbb{I}_{d}}=d$ and hence, these copulas are also “nearly independent” (see paragraph below (1.2)); several popular Archimedean copulas models as, e.g., Frank copula, Clayton copula and Ali-Mikhail-Haq copula (cf. [4, Table 1]) fall in this class exhibiting both pairwise and mutual asymptotic independence. In contrast, there are as well Archimedean copulas exhibiting only pairwise asymptotic independence but not mutual asymptotic independence. The following result provides sufficient conditions on the generator $\phi$ to obtain such Archimedean copulas.
@@
Let the dependence structure of a random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ be given by an Archimedean copula $C^{\phi}$ with generator $\phi$ as in (2.8). Suppose $\phi^{\prime}(1)=0$ and
@@
is a positive function, which is slowly varying at 00. Then $\boldsymbol{Z}$ possesses pairwise asymptotic independence but does not possess mutual asymptotic independence.
@@
where $\ell$ is a slowly varying function at 0 (follows from [4, Corollary 4.7]). Hence, the upper tail order of $C_{S}$ is $\kappa_{S}=1$ for all $S\subset\mathbb{I}_{d}$ with $|S|\geq 2$. To obtain an example of such a copula, take some parameter $\theta\in(0,\infty)$ and define the generator
@@
Let the dependence structure of a random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ be given by a Gaussian copula $C^{\Sigma}$ with positive-definite correlation matrix $\Sigma$. Then $\boldsymbol{Z}$ exhibits mutual asymptotic independence if and only if $\Sigma^{-1}_{S}\boldsymbol{1}_{|S|}>\boldsymbol{0}_{|S|}$ for all non-empty sets $S\subseteq\mathbb{I}_{d}$.
The proof is quite involved, requiring a few auxiliary results based on the recently derived knowledge ([12, 11]) on the asymptotic behavior of tail probabilities of a multivariate distribution with identically Pareto margins and Gaussian copula $C^{\Sigma}$. Curiously, the ingredients of the proof allow us to find the tail asymptotics of the survival copula of any $|S|$-dimensional marginal in terms of its tail order.
@@
A few interesting features of Proposition 3.10 and related results (see Appendix A) are to be noted here.
  1. 1.
    Although Proposition 3.10 only gives the tail order of $\widehat{C}^{\Sigma}_{S}$, in fact, the exact tail asymptotics for $\widehat{C}^{\Sigma}_{S}$ is available in Proposition A.3.
  2. 2.
    The upper tail order $\kappa_{S}$ in (3.2) is obtained as a solution to a quadratic programming problem as given in Lemma A.1.
  3. 3.
    With respect to (3.2), for subsets $S,T\subset\mathbb{I}_{d}$ with $S\subsetneq T$ and $|S|\geq 2$, it is possible to have (i) $\kappa_{S}<\kappa_{T}$, (ii) $\kappa_{S}=\kappa_{T}$, with $\ell_{S}(u)\sim c\,\ell_{T}(u),u\downarrow 0$ for $c>0$, and (iii) $\kappa_{S}=\kappa_{T}$, with $\ell_{S}(u)=o\left(\ell_{T}(u)\right),u\downarrow 0$. In Example 3.12, we can observe both (i) and (ii) holding under different assumptions; an example for (iii) with Pareto margins and Gaussian copula is available in [11, Remark 5].
@@
For the purpose of illustration, we provide a positive-definite correlation matrix (with $d=3$) for a Gaussian copula parametrized by a single parameter $\rho$ which exhibits mutual asymptotic independence for only certain values of $\rho$ and only pairwise asymptotic independence for other feasible values; see Das and Fasen-Hartmann [11, Ex. 1(b)] for further details. Throughout, we denote by $\ell_{j}(u),j\in\mathbb{N}$, a slowly varying function at 0. Consider the Gaussian copula $C^{\Sigma}$ with correlation matrix
@@
    Suppose $\rho<1/({2\sqrt{2}-1})\approx 0.55$. Then one can check that $\Sigma^{-1}\boldsymbol{1}>\boldsymbol{0}$, and hence, mutual asymptotic independence holds as well. In fact, we can find the behavior of the survival copula (using Proposition A.3 or, Das and Fasen-Hartmann [11, Example 1(b)]) as follows: as $u\downarrow 0$,
@@
    $\widehat{C}_{13}^{\Sigma}(u,u)$ $=\widehat{C}_{23}^{\Sigma}(u,u)\sim u^{\frac{2}{1+\sqrt{2}\rho}}\ell_{2}(u),\quad\text{and,}$, (3.4)
    $\widehat{C}_{12}^{\Sigma}(u,u)$ $\sim u^{\frac{2}{1+\rho}}\ell_{3}(u),$ (3.5)
@@
    On the other hand, if $\rho\geq 1/({2\sqrt{2}-1})$, then $\Sigma^{-1}\boldsymbol{1}\ngtr\boldsymbol{0}$ and the copula does not have mutual asymptotic independence. Note that in this case, the behavior of the two-dimensional marginal survival copulas will still be given by (3.4) and (3.5), but the tail behavior as seen in (3.3) does not hold anymore. Now, as $u\downarrow 0$, we have
@@
    In fact we can check that $\ell_{4}(u)\sim\beta\,\ell_{3}(u)$ for some constant $\beta>0$ (cf. [11, Example 1(b)])), and hence
    $\widehat{C}^{\Sigma}(u,u,u)\sim\beta\,\widehat{C}_{12}^{\Sigma}(u,u),\quad u\downarrow 0,$

### v3

@@ 3 Mutual asymptotic independence
Pairwise asymptotic independence has often either been used as a natural extension of asymptotic independence [2, 19], or taken as a consequence from other relevant properties de Haan and Ferreira [13, Remark 6.2.5], or implicitly assumed [27] in a variety of works. Next, we define a notion that captures the global joint concurrent tail behavior of random vectors portrayed by many popular multivariate dependence structures, e.g., dependence defined using Gaussian, Marshall-Olkin, or various Archimedean copulas, etc., but not restricted to mere replication of pairwise comparisons of tails.
@@
A random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with continuous marginal distributions and survival copula $\widehat{C}$ is mutually asymptotically independent if for all $S\subseteq\mathbb{I}_{d}$ with $|S|\geq 2$, we have
@@
Remark 3.2.
Some explanation is due here in order to distinguish between the traditional notion of multivariate asymptotic independence (also called “mutual” asymptotic independence sometimes) in dimensions $d>2$ (Resnick [40, Chapter 5.5], Galambos [17, Chapter 5.2], McNeil, Frey and Embrechts [34, Chapter 7.6]) and the notion defined in Definition 3.1. Due to Resnick [40, Proposition 5.27] under the constrain that the margins are continuous, identically distributed and in the maximum domain of attraction of a univariate extreme value distribution, pairwise asymptotic independence is equivalent to a distribution having “multivariate asymptotic independence” meaning that the distribution lies in the maximum domain of attraction of an extreme value distribution with independent marginals (the limit distribution is a product measure). But our notions of “mutual asymptotic independence” and “multivariate asymptotic independence” are not equivalent which we see in the following and in particular, in Example 3.4.
When $d=2$, both (2.2) and (3.1) boil down to (2.1) and hence, are equivalent. Assuming $d\geq 3$ and mutual asymptotic independence, if we take all choices of $S\subseteq\mathbb{I}_{d}$ with $|S|=2$, then (3.1) is just a restatement of (2.2), implying pairwise asymptotic independence. We summarize this in the next proposition.
@@
If a random vector $\boldsymbol{Z}\in\mathbb{R}^{d},$ $d\geq 2$ with continuous marginal distributions is mutually asymptotically independent, then it is also pairwise asymptotically independent.
The reverse implication of Proposition 3.3 is not necessarily true as we see in the following example, which mimics the consequences for the analogous notions of classical “mutual” and “pairwise independence” ([23]).
@@
The difference between pairwise and mutual independence can be shown using an $\mathbb{R}^{3}$-valued random vector with Bernoulli marginals (cf. Hogg, McKean and Craig [23, Chapter 2]). We take a similar spirit using uniform marginals. Consider i.i.d. uniform [0,1] random variables $U,V,W$. Then ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{Z}^{*}}=(U,V,W)$ is mutually asymptotically independent (cf. Example 3.5) and hence, pairwise asymptotically independent as well. Now consider $\boldsymbol{Z}=(Z_{1},Z_{2},Z_{3})\sim F$ such that
[…]
First note that for $0<z<1$, marginally,
@@
and hence, the $Z_{j}$’s are identically distributed.
  • (i)
    If $\widehat{C}$ denotes the survival copula of $\boldsymbol{Z}$, then we can check that for any $\{j,\ell\}\subset\{1,2,3\}$,
[…]
@@
    $=(\sqrt{1+3u}-1)^{2}=9u^{2}/4+o(u^{2}),\quad u\downarrow 0.$ (3.2)
    Hence, $\boldsymbol{Z}$ exhibits pairwise asymptotic independence.
  • (ii)
    But
@@
  • (iii)
    We can compute that if $\boldsymbol{Z}^{(1)},\ldots,\boldsymbol{Z}^{(n)}$ are i.i.d. $F$ and $\boldsymbol{M}_{n}$ is the random vector of component-wise maxima given by
    $\boldsymbol{M}_{n}=\left(\bigvee_{i=1}^{n}Z_{1}^{(i)},\bigvee_{i=1}^{n}Z_{2}^{(i)},\bigvee_{i=1}^{n}Z_{3}^{(i)}\right),$
    then with $a_{n}=3/{2n}$ and $b_{n}=1$, we have for $\boldsymbol{x}\in\mathbb{R}$,
    $\lim_{n\to\infty}\mathbb{P}(\boldsymbol{M}_{n}\leq a_{n}\boldsymbol{x}+b_{n})=\lim_{n\to\infty}F^{n}(a_{n}\boldsymbol{x}+b_{n})=G(\boldsymbol{x})=\prod_{i=1}^{3}\Psi_{1}(x_{i}).$ (3.3)
    where $\Psi_{1}(\cdot)$ is a univariate extreme value distribution given by $\Psi_{1}(x)=\min(e^{x},1),x\in\mathbb{R}$. Thus $F\in\mathrm{MDA}(G)$ where $G$ is indeed a product measure according to (3.3), implying multivariate asymptotic independence although $F$ does not have mutual asymptotic independence as shown in (ii).
    For illustration, we showed the multivariate asymptotic independence by hand, but the pairwise asymptotic independence in (i) and Resnick [40, Proposition 5.27] already imply multivariate asymptotic independence.
@@ 3.1 Examples: Part I
In Example 2.4 we stated that any random vector $\boldsymbol{Z}$ with dependence given by the generalized Marshall-Olkin survival copula $\widehat{C}^{\mathrm{MO}(\Lambda)}$ as defined in (2.5) is pairwise asymptotically independent. In fact, Lin and Li [32, Proposition 3] allows us to conclude that $\widehat{C}^{\mathrm{MO}(\Lambda)}$ is indeed mutually asymptotically independent as well.
@@ 3.2 Examples: Part II
In this section, we discuss examples that are pairwise asymptotically independent but sometimes are not mutually asymptotically independent. This will include a large class of examples from the Archimedean copula and the Gaussian copula family.
3.2.1 Archimedean copula
@@
Let the dependence structure of a random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with continuous marginal distributions be given by an Archimedean copula $C^{\phi}$ with generator $\phi$ as in (2.8). Suppose $\phi^{\leftarrow}$ is $d$-times continuously differentiable and $(-D)^{j}\phi^{\leftarrow}(0)<\infty$ $\forall\,j\in\mathbb{I}_{d}$. Then $\boldsymbol{Z}$ possesses both pairwise and mutual asymptotic independence.
The proof follows directly from Charpentier and Segers [4, Theorem 4.3]. The Archimedean copulas of Theorem 3.7 have the property that for any subset $S\subset\mathbb{I}_{d}$ with $|S|\geq 2$, the survival copula $\widehat{C}_{S}^{\phi}$ of the $|S|$-dimensional marginal behaves like the independence copula near the tails, i.e.,
@@
where the upper tail order of $C_{S}$ is $\kappa_{S}=|S|$ (also follows from [4, Theorem 4.3]). In particular, the upper tail order for $C^{\phi}$ is $\kappa=\kappa_{\mathbb{I}_{d}}=d$ and hence, these copulas are also “nearly independent” (see paragraph below (1.2)); several popular Archimedean copulas models, e.g. as Frank copula, Clayton copula and Ali-Mikhail-Haq copula (Charpentier and Segers [4, Table 1]) fall in this class exhibiting both pairwise and mutual asymptotic independence. In contrast, there are also Archimedean copulas exhibiting only pairwise asymptotic independence but not mutual asymptotic independence. The following result provides sufficient conditions on the generator $\phi$ to obtain such Archimedean copulas.
@@
Let the dependence structure of a random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with continuous marginal distributions be given by an Archimedean copula $C^{\phi}$ with generator $\phi$ as in (2.8). Suppose $\phi^{\prime}(1)=0$ and
@@
is a positive function, which is slowly varying at 0. Then $\boldsymbol{Z}$ possesses pairwise asymptotic independence but does not possess mutual asymptotic independence.
@@
where $\ell$ is a slowly varying function at 0 (follows from Charpentier and Segers [4, Corollary 4.7]). Hence, the upper tail order of $C_{S}$ is $\kappa_{S}=1$ for all $S\subset\mathbb{I}_{d}$ with $|S|\geq 2$. To obtain an example of such a copula, take some parameter $\theta\in(0,\infty)$ and define the generator
@@
Let the dependence structure of a random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with continuous marginal distributions be given by a Gaussian copula $C^{\Sigma}$ with positive-definite correlation matrix $\Sigma$. Then $\boldsymbol{Z}$ exhibits mutual asymptotic independence if and only if $\Sigma^{-1}_{S}\boldsymbol{1}_{|S|}>\boldsymbol{0}_{|S|}$ for all non-empty sets $S\subseteq\mathbb{I}_{d}$.
The proof of the theorem is quite involved, requiring a few auxiliary results based on the recently derived knowledge on the asymptotic behavior of tail probabilities of a multivariate distribution with identically Pareto marginals and Gaussian copula $C^{\Sigma}$ in Das and Fasen-Hartmann [11]. Hence the proof has been relegated to Appendix A. Curiously, the ingredients of the proof allow us to find the tail asymptotics of the survival copula of any $|S|$-dimensional marginal in terms of its tail order.
@@
A proof of this result is given in Appendix A as well.
@@
A few interesting features of Proposition 3.10 and related results are to be noted here.
  1. (a)
    Although Proposition 3.10 only gives the tail order of $\widehat{C}^{\Sigma}_{S}$, in fact, the exact tail asymptotics for $\widehat{C}^{\Sigma}_{S}(u\boldsymbol{v}_{S})$ as $u\downarrow 0$ for $\boldsymbol{v}_{S}=(v_{s})_{s\in S}$, $v_{s}\in(0,1)$, including the slowly varying function is available in Proposition A.3 in Appendix A.
  2. (b)
    The upper tail order $\kappa_{S}$ in (3.4) is obtained as a solution to a quadratic programming problem; the exact solution is given in Lemma A.1 in Appendix A.
  3. (c)
    With respect to (3.4), for subsets $S,T\subset\mathbb{I}_{d}$ with $S\subsetneq T$ and $|S|\geq 2$, it is possible to have (i) $\kappa_{S}<\kappa_{T}$, (ii) $\kappa_{S}=\kappa_{T}$, with $\ell_{S}(u)\sim c\,\ell_{T}(u),u\downarrow 0$ for $c>0$, and (iii) $\kappa_{S}=\kappa_{T}$, with $\ell_{S}(u)=o\left(\ell_{T}(u)\right),u\downarrow 0$. In Example 3.12, we can observe both (i) and (ii) holding under different assumptions; an example for (iii) with Pareto marginals and Gaussian copula is available in Das and Fasen-Hartmann [11, Remark 5].
  4. (d)
    In Hua and Joe [24, Example 1], the authors already state that the tail order $\kappa$ of a Gaussian copula with positive definite correlation matrix $\Sigma$ is $\kappa=\boldsymbol{1}^{\top}\Sigma^{-1}\boldsymbol{1}$ (cf. Joe [26, Section 4.3.2]). However, to the best of our knowledge, the aforementioned paper does not specify that $\Sigma^{-1}\boldsymbol{1}>\boldsymbol{0}$ is indeed a necessary condition for the result, since otherwise the statement is not valid; in fact, if $\Sigma^{-1}\boldsymbol{1}\ngtr\boldsymbol{0}$ then $\kappa<\boldsymbol{1}^{\top}\Sigma^{-1}\boldsymbol{1}$ is possible (cf. Lemma A.1).
@@
For the purpose of illustration, we provide a positive-definite correlation matrix (with $d=3$) for a Gaussian copula parametrized by a single parameter $\rho$ which exhibits mutual asymptotic independence for only certain values of $\rho$ and only pairwise asymptotic independence for other feasible values; see Das and Fasen-Hartmann [11, Example 1(b)] for further details. Throughout, we denote by $\ell_{j}(u),j\in\mathbb{N}$, a slowly varying function at 0. Consider the Gaussian copula $C^{\Sigma}$ with correlation matrix
@@
    Suppose $\rho<1/({2\sqrt{2}-1})\approx 0.55$. Then one can check that $\Sigma^{-1}\boldsymbol{1}>\boldsymbol{0}$, and hence, mutual asymptotic independence holds as well. In fact, we can find the behavior of the survival copula (using Proposition A.3 or, Das and Fasen-Hartmann [11, Example 1(b)]): As $u\downarrow 0$,
@@
[…]
@@
    On the other hand, if $\rho\geq 1/({2\sqrt{2}-1})$, then $\Sigma^{-1}\boldsymbol{1}\ngtr\boldsymbol{0}$ and the copula does not have mutual asymptotic independence. Note that in this case, the behavior of the two-dimensional marginal survival copulas will still be given by (3.6), but the tail behavior as seen in (3.5) does not hold anymore. Now, as $u\downarrow 0$, we have
@@
    In fact we can check that $\ell_{4}(u)\sim\beta\,\ell_{3}(u)$ as $u\downarrow 0$ for some constant $\beta>0$ (Das and Fasen-Hartmann [11, Example 1(b)]), and hence
    $\widehat{C}^{\Sigma}(u,u,u)\sim\beta\,\widehat{C}_{\{12\}}^{\Sigma}(u,u),\quad u\downarrow 0,$